# Structure of the free interfaces near triple junction singularities in harmonic maps and optimal partition problems

### v1

@@ Abstract.
We consider energy-minimizing harmonic maps into trees and we prove the regularity of the singular part of the free interface near triple junction points. Precisely, by a dimension reduction and a topological argument, we prove that there are no admissible frequencies between $1$ and $\nicefrac{{3}}{{2}}$, and we classify the $\nicefrac{{3}}{{2}}$-homogeneous solutions. Then, by proving a new epiperimetric inequality, we show that around any point of frequency $\nicefrac{{3}}{{2}}$, the free interface is composed of three $C^{1,\alpha}$-smooth $(d-1)$-dimensional manifolds (of points of frequency $1$) with common $C^{1,\alpha}$-regular boundary (made of points of frequency $\nicefrac{{3}}{{2}}$) that meet along this boundary at 120 degree angles, which in particular implies that the free interface is a smooth deformation of the $Y$ cone. Our results also apply to spectral optimal partition problems for the Dirichlet eigenvalues. Moreover, as a byproduct, we also prove that the $Y$-configuration is the only minimizer for the min-max optimal $3$-partition problem for the Dirichlet eigenvalues on the sphere, in any dimension.
@@ 1. Introduction
(1.1) $\Sigma_{N}:=\left\{X\in\mathbb{R}^{N}\colon X_{i}X_{j}=0~{}\text{for all }j\neq i\right\}\,,$
By the works of Gromov-Schoen [GS92], Caffarelli-Lin [CL07, CL08], Conti-Terracini-Verzini [CTV02, CTV03, CTV05a] it is known that if
@@
then $\mathcal{F}(u)$ can be decomposed into regular and singular parts according to the values of the Almgren’s frequency $\gamma(u,x)$ as follows:
$\mathrm{Reg}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=1\big{\}}\qquad\text{and}\qquad\mathrm{Sing}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)>1\big{\}}.$
Gromov-Schoen and Caffarelli-Lin [GS92, CL08] studied the regular part $\mathrm{Reg}(u)$ in every dimension and showed that it is a smooth $(d-1)$-dimensional manifold, while the structure of the singular set was studied in [GS92] and [Alp20, Dee22], where it was shown that $\mathrm{Sing}(u)$ has Hausdorff dimension $d-2$ ([GS92]) and is a $(d-2)$-rectifiable ([Alp20, Dee22]). Moreover, Conti, Terracini and Verzini showed in [CTV03] that the singular set in dimension $2$ is discrete and gave a complete description of the nodal set $\mathcal{F}(u)$ around any singular point.
In this paper we study the regularity of the singular set $\mathrm{Sing}(u)$ (and the structure of $\mathcal{F}(u)$ around singular points) in any dimension. Our main result is the following:
@@
Let $u:B_{1}\to\Sigma_{N}$ be an energy-minimizing map in $B_{1}\subset\mathbb{R}^{d}$, for some $d\geq 2$ and $N\geq 1$. Then, the singular set can be decomposed as the disjoint union of
$\mathrm{Sing}(u)=\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)>\nicefrac{{3}}{{2}}\Big{\}}.$
@@
The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u):=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold. Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, the nodal set $\mathcal{F}(u)$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$.
The main ingredients for the proof of Theorem 1.1 are the classification of the singular homogeneous solutions of lowest frequency, which we achieve in Theorem 1.2 via a topological lemma combined with a dimension reduction argument (see Section 2) and a new epiperimetric inequality (Theorem 1.3) which we obtain via a contradiction argument in the spirit of Weiss [Wei99] (see also [GPS16, FS16]). Furthermore, Theorem 1.1 solves the problem of finding the optimal frequency gap in every dimension $d\geq 2$. Using this result, in Theorem 1.5 we also solve, in any dimension, an open problem by Helffer, Hoffmann-Ostenhof and Terracini in [HHT09], which is the limit case of the Bishop-Friedland-Hayman’s conjecture [Bis92, FH76]. Finally, in Theorem 1.6 and Theorem 1.7, we discuss two corollaries of Theorem 1.1 in the context of harmonic maps and optimal partition problems.
@@
Results in the spirit of Theorem 1.1 exist in the context of area-minimizing currents, where the nodal set $\mathcal{F}(u)$ is replaced by the support of an area-minimizing current (mod $p=3$); in this framework, description of the triple junction singularities is known thanks to the historical works from the 70s of Jean Taylor [Tay73, Tay76], for hypersurfaces in $\mathbb{R}^{3}$, and from the early 90s of Leon Simon [Sim93], in any dimension $d\geq 3$ (see also the lecture notes [Min24]); we also remark that the analysis for general co-dimension 1 area-minimizing currents mod p was completed only recently in [DHMS20], [DHM+21] and [DHM+22]. To our knowledge Theorem 1.1 is the first instance of a study of triple junction singularities in the contexts of free boundary problems and harmonic maps in dimension $d>2$; we also notice that in the framework of Theorem 1.1 the classification of the tangent maps was (until now) an open problem even in dimension $d=3$ and cannot be obtained via elementary geometric arguments. The epiperimetric inequality approach to the analysis of the singular points in the context of harmonic maps is, to our knowledge, also new in this field.
In the rest of the introduction we give a detailed account on the known results (Section 1.1); in Section 1.2 we state our optimal gap result (Theorem 1.2) and the epiperimetric inequality (Theorem 1.3); finally, Section 1.3 is dedicated to the Helffer-Hoffmann-Ostenhof-Terracini’s problem and the Bishop-Friedland-Hayman’s conjecture.
1.1. Notations and known results
Let us fix the dimension $d\geq 2$. We work with a general open set $D\subseteq\mathbb{R}^{d}$ and with functions $u=(u_{1},\dots,u_{N})\colon D\to\mathbb{R}^{N}$ with $N\geq 2$ components whose support in $D$ is disjoint, i.e. $u_{i}u_{j}\equiv 0$ in $D$. More precisely, we consider the singular (non-positively curved) target space
$\Sigma_{N}:=\Bigg{\{}X\in\mathbb{R}^{N}\colon\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{N}X_{i}^{2}X_{j}^{2}=0\Bigg{\}}$ =$=\left\{X\in\mathbb{R}^{N}\colon X_{i}X_{j}=0~{}\text{for all }j\neq i\right\}\,,$
on which it is naturally defined the distance
We then consider the Sobolev-type space
$H^{1}(D;\Sigma_{N})=\left\{u\in(H^{1}(D))^{N}\colon u_{i}u_{j}=0~{}\text{a.e. in }D~{}\text{for all }i\neq j\right\}.$
@@ 1.1. Notations and known results
The space $H^{1}_{\textup{loc}}(D;\Sigma_{N})$ is analogously defined. Now, to each open $D\subseteq\mathbb{R}^{d}$ and each function $u\in H^{1}(D;\Sigma_{N})$ we associate the corresponding Dirichlet energy, defined as
@@
for all open $\Omega\Subset D$ and all $v\in H^{1}(\Omega;\Sigma_{N})$ such that $u-v\in H^{1}_{0}(\Omega;\Sigma_{N})$. On the other hand, some of our results hold true in a larger class of “critical points”, namely, the so called class $\mathcal{S}$. In order to define it, we first introduce some notation which is also useful in the rest of the paper. Given $u\in\mathcal{A}(D;N)$ and a function
$a\colon\{1,\dots,N\}\to\{-1,1\}$
we define the real-valued function $a\cdot u\in H^{1}_{\textup{loc}}(D)$ as
$(a\cdot u)(x):=\sum_{i=1}^{N}a(i)u_{i}(x)\quad\text{for }x\in D.$
Moreover, for any $i\in\{1,\dots,N\}$, we denote by $\sigma_{i}:=\{1,\dots,N\}\to\{-1,1\}$ the function
$\sigma_{i}(k):=\begin{cases}1,&\text{if }k=i,\\ -1,&\text{if }k\neq 1,\end{cases}$
so that
$\sigma_{i}\cdot u=u_{i}-\sum_{j\neq i}u_{j}\quad\text{in $D$, for all }i=1,\dots,N.$
Now, we say that a function $u\in\mathcal{A}(D;N)$ belongs to the class $\mathcal{S}(D;N)$ if
$-\Delta u_{i}&\leq 0,&&\text{in }D,\\ -\Delta(\sigma_{i}\cdot u)&\geq 0,&&\text{in }D,\end{aligned}\right.$
in the sense of distributions, for all $i=1,\dots,N$. In particular, there holds $\mathcal{M}(D;N)\subseteq\mathcal{S}(D;N)$ (see e.g. [CTV05c, Theorem 5.1]).
The aim of present paper is to investigate regularity questions for solutions $u$ in the classes $\mathcal{S}(D;N)$ and $\mathcal{M}(D;N)$ focusing, in particular, on the geometrical structure and regularity of the free boundary, defined as
$\mathcal{F}(u):=\bigcup_{i=1}^{N}\partial\{x\in D\colon u_{i}(x)>0\},$
in the interior of $D$. We point out that $u\in C^{0,1}_{\textup{loc}}(D;\mathbb{R}^{N})$, see e.g. [CTV05c], which implies that the sets

### v2

@@ Abstract.
We consider energy-minimizing harmonic maps into trees and we prove the regularity of the singular part of the free interface near triple junction points. Precisely, by proving a new epiperimetric inequality, we show that around any point of frequency $\nicefrac{{3}}{{2}}$, the free interface is composed of three $C^{1,\alpha}$-smooth $(d-1)$-dimensional manifolds (composed of points of frequency $1$) with common $C^{1,\alpha}$-regular boundary (made of points of frequency $\nicefrac{{3}}{{2}}$) that meet along this boundary at 120 degree angles. Our results also apply to spectral optimal partition problems for the Dirichlet eigenvalues.
@@ 1. Introduction
(1.1) $\Sigma_{N}:=\left\{X\in\mathbb{R}^{N}\colon X_{i}X_{j}=0~{}\text{for all }j\neq i\right\}\,$
endowed with the distance
By the works of Gromov-Schoen [GS92], Caffarelli-Lin [CL07, CL08], Tavares-Terracini [TT12] and Soave-Terracini [ST15] it is known that if
@@
then $\mathcal{F}(u)$ can be decomposed into regular and singular parts according to the values of the Almgren’s frequency $\gamma(u,x)$ (see Section 2.2) as follows:
$\mathrm{Reg}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=1\big{\}}\qquad\text{and}\qquad\mathrm{Sing}(u)=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)\geq\nicefrac{{3}}{{2}}\big{\}}.$
By [GS92, CL08], in every dimension $d\geq 2$, the regular part $\mathrm{Reg}(u)$ is a smooth $(d-1)$-dimensional manifold, while the structure of the singular set was studied in [GS92] and [Alp20, Dee22], where it was shown that $\mathrm{Sing}(u)$ has Hausdorff dimension $d-2$ ([GS92]) and is $(d-2)$-rectifiable ([Alp20, Dee22]). In dimension $d=2$, Conti, Terracini and Verzini showed in [CTV03] that the singular set is discrete and gave a complete description of the nodal set $\mathcal{F}(u)$ around any singular point (see also the recent paper [LMNS24] for the case of generic solutions).
In this paper we study the regularity of the lowest stratum of $\mathrm{Sing}(u)$, that is
(1.2) $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u):=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\big{\}},$
and the structure of the entire free interface $\mathcal{F}(u)$ around singular points in $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ in any dimension $d\geq 2$. Our main result is the following:
@@
Let $u:B_{1}\to\Sigma_{N}$ be an energy-minimizing map in $B_{1}\subset\mathbb{R}^{d}$, for some $d\geq 2$ and $N\geq 1$. Then, the singular set of $\mathrm{Sing}{(u)}$ of the free interface $\mathcal{F}(u)$ can be decomposed as
$\mathrm{Sing}{(u)}=\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)\cup\Big{\{}x\in\mathcal{F}{(u)}\ :\ \gamma(u,x)\geq\nicefrac{{3}}{{2}}+\delta_{d}\Big{\}},$
@@
where $\delta_{d}>0$ is dimensional constant. The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ is an open subset of $\mathrm{Sing}(u)$ and, locally, a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold. Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, the nodal set $\mathcal{F}(u)$ consists of three $(d-1)$-manifolds $\Gamma_{12}$, $\Gamma_{23}$, $\Gamma_{31}$ with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}(u)$.
@@
Results in the spirit of Theorem 1.1 exist in the context of area-minimizing currents, where the nodal set $\mathcal{F}(u)$ is replaced by the support of an area-minimizing current (mod $p=3$); in this framework, description of the triple junction singularities is known thanks to the historical works from the 70s of Jean Taylor [Tay73, Tay76], for hypersurfaces in $\mathbb{R}^{3}$, and from the early 90s of Leon Simon [Sim93], in any dimension $d\geq 3$ (see also the lecture notes [Min24]); we also remark that the analysis for general co-dimension 1 area-minimizing currents mod p was completed only recently in [DHMS20], [DHM+21] and [DHM+22]. To our knowledge Theorem 1.1 is the first instance of a study of triple junction singularities in the contexts of free boundary problems and harmonic maps in dimension $d>2$.
The main ingredient in the proof of Theorem 1.1 is the following epiperimetric inequality (Theorem 1.2) for the $\nicefrac{{3}}{{2}}$-Weiss energy
$W_{\frac{3}{2}}(u):=\sum_{i=1}^{N}\int_{B_{1}}|\nabla u_{i}|^{2}\,\mathrm{d}x-\frac{3}{2}\sum_{i=1}^{N}\int_{\partial B_{1}}u_{i}^{2}\,\mathrm{d}S$
near homogeneous triple junctions $Y(x)=Y(x_{d-1},x_{d})$ with $2$-dimensional profile defined as
(1.3) $Y_{i}(r,\theta)=\begin{cases}r^{\frac{3}{2}}\left|\cos\left(\frac{3}{2}\theta\right)\right|,&\text{for }\pi-\frac{2\pi}{3}(i-1)\leq\theta\leq\pi-\frac{2\pi}{3}i,\\ 0,&\text{elsewhere},\end{cases}$
for $i=1,2,3$.
Theorem 1.2 (Epiperimetric inequality).
and
$W_{\frac{3}{2}}(u)\leq(1-\varepsilon)W_{\frac{3}{2}}(c).$
The proof of Theorem 1.2 is obtained via a contradiction argument in the spirit of Weiss [Wei99] (see also [FS16, GPS16]). The epiperimetric inequality approach to the analysis of the singular points in the context of harmonic maps is, to our knowledge, new in this field.
1.1. Two consequences of Theorem 1.1
We now discuss two corollaries of Theorem 1.1 for harmonic maps with values in metric graphs and for spectral optimal partitions.
First of all we notice that an immediate corollary of Theorem 1.1 is that the same result holds for harmonic maps with values in a locally finite tree $\Sigma$. Indeed, by [GS92, Dee22], if $u:B_{1}\to\Sigma$ is an energy-minimizing map defined in $B_{1}\subset\mathbb{R}^{d}$ with values in a locally finite metric tree, and if $p$ is a vertex of $\Sigma$, then the set $\mathcal{F}{(u)}:=u^{-1}(\{p\})$ can be decomposed into a regular and a singular parts, $\mathrm{Reg}{(u)}$ and $\mathrm{Sing}{(u)}$, where $\mathrm{Reg}{(u)}$ is a a smooth $(d-1)$-manifold that consists of the points of order (frequency) ${\rm Ord}^{u}=1$, while $\mathrm{Sing}{(u)}$ is $(d-2)$-rectifiable and consists of of points of order (frequency) ${\rm Ord}^{u}>1$; furthermore, for every $x\in\mathcal{F}{(u)}$, there is a neighborhood $\Omega\subset B_{1}$ such that $u:\Omega\to\Sigma$ has values in a subtree $\Sigma_{N}$ of the form (1.1). Thus, we have the following.
Theorem 1.3.
Let $\Sigma$ be a metric tree which is locally finite and let $u:B_{1}\to\Sigma$ be an energy-minimizing map in $B_{1}\subset\mathbb{R}^{d}$, for some $d\geq 2$. Then, for every vertex $p$ of $\Sigma$ the singular set of $\mathcal{F}{(u)}:=u^{-1}(\{p\})$ can be decomposed as
$\mathrm{Sing}{(u)}=\Big{\{}x\in\mathcal{F}(u)\ :\ {\rm Ord}^{u}(x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}(u)\ :\ {\rm Ord}^{u}(x)\geq\nicefrac{{3}}{{2}}+\delta_{d}\Big{\}},$
where $\delta_{d}>0$ is a dimensional constant. The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}:=\big{\{}x\in\mathcal{F}{(u)}\ :\ {\rm Ord}^{u}(x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold. Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$, the set $\mathcal{F}{(u)}$ consists of three $(d-1)$-manifolds with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$.
Another consequence concerns the regularity of the following optimal partition problem studied in [CL07], [CTV03, CTV05a] and [OV24]. Given a bounded open set $D\subset\mathbb{R}^{d}$, $d\geq 2$, and a natural number $N\geq 2$ we consider the optimal partition problem
$\mathcal{F}{(u)}:=D\cap\Big{(}{\bigcup_{j=1}^{N}}\partial\Omega_{j}\Big{)}$
can be decomposed as a regular part and a singular part as follows (see [CL07, ST15, Alp20]): the regular part $\mathrm{Reg}{(u)}$ consists of all points $x\in\mathcal{F}{(u)}$ of frequency $\gamma(u,x)=1$ and is a smooth manifold of dimension $d-1$; the singular part $\mathrm{Sing}{(u)}$ is a closed $(d-2)$-rectifiable set composed of all the points of frequency $\gamma(u,x)\geq\nicefrac{{3}}{{2}}$. We have the following theorem.
Theorem 1.4.
Let $(\Omega_{1},\dots,\Omega_{N})$ be a minimizer of (1.4) and $u=(u_{1},\dots,u_{N})$ be the corresponding vector of the normalized first eigenfunctions. Then, the singular set of $\mathrm{Sing}{(u)}$ of the free interface $\mathcal{F}(u)$ can be decomposed as
$\mathrm{Sing}{(u)}=\Big{\{}x\in\mathcal{F}{(u)}\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\Big{\}}\cup\Big{\{}x\in\mathcal{F}{(u)}\ :\ \gamma(u,x)\geq\nicefrac{{3}}{{2}}+\delta_{d}\Big{\}},$
where $\delta_{d}>0$ is dimensional constant. The set $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}:=\big{\{}x\in\mathcal{F}(u)\ :\ \gamma(u,x)=\nicefrac{{3}}{{2}}\big{\}}$ is an open subset of $\mathrm{Sing}(u)$ and is locally a $(d-2)$-dimensional $C^{1,\alpha}$-smooth manifold. Moreover, around any $x\in\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$, the set $\mathcal{F}{(u)}$ consists of three $(d-1)$-manifolds with common boundary $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$, which are all $C^{1,\alpha}$-regular up to $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$ and form $120$ degree angles at $\mathcal{F}_{\nicefrac{{3}}{{2}}}{(u)}$.
The rest of the paper is organized as follows. In Section 2 we fix the notation, the functional setting and we recall some known facts about segregated solutions. In Section 3, we classify the $\nicefrac{{3}}{{2}}$-homogeneous solutions and we find, as a corollary, the shape of minimizers of a certain optimal partition problem on the sphere (see Corollary 3.6). Section 5 is devoted to the proof of our main tool, the epiperimetric inequality Theorem 1.2 and, finally, in Section 6 we conclude the proof of our main regularity result Theorem 1.1.
2. Preliminaries
In this section we introduce the functional setting and we define the main tools in our framework, such as monotonicity formulas.
For any $N\geq 1$, let $\Sigma_{N}$ be as in (1.1). For any open set $D\subseteq\mathbb{R}^{d}$, we define the Sobolev space
$H^{1}(D;\Sigma_{N})=\left\{u\in(H^{1}(D))^{N}\colon u_{i}u_{j}=0~{}\text{a.e. in }D~{}\text{for all }i\neq j\right\},$
@@ 2. Preliminaries
with the space $H^{1}_{\textup{loc}}(D;\Sigma_{N})$ being analogously defined; we refer to [KS93] for the theory of Sobolev spaces in this setting. Now, to each open $D\subseteq\mathbb{R}^{d}$ and each function $u\in H^{1}(D;\Sigma_{N})$ we associate the corresponding Dirichlet energy, defined as
@@
for all open $\Omega\Subset D$ and all $v\in H^{1}(\Omega;\Sigma_{N})$ such that $u-v\in H^{1}_{0}(\Omega;\Sigma_{N})$. We also introduce the class $\mathcal{S}(D;N)$ of functions $u\in\mathcal{A}(D;N)$ satisfying
$-\Delta u_{i}&\leq 0,&&\text{in }D,\\ -\Delta\Big{(}u_{i}-\sum_{j\neq i}u_{j}\Big{)}&\geq 0,&&\text{in }D,\end{aligned}\right.$
in the sense of distributions, for all $i=1,\dots,N$. In particular, there holds $\mathcal{M}(D;N)\subseteq\mathcal{S}(D;N)$ (see e.g. [CTV05b, Theorem 5.1]).
We now recall the main basic properties of critical points $u\in\mathcal{S}(D;N)$ and energy-minimizers $u\in\mathcal{M}(D;N)$ and of the free boundary $\mathcal{F}(u)$. Most of the properties we recall were proved in a series of papers, namely [GS92, CL07] for the class $\mathcal{M}(D;N)$ and [CTV05b, TT12, ST15] for $\mathcal{S}(D;N)$; for what concerns the regularity of different classes of segregated solutions we refer to [NRS24, ST24b, ST24a].
2.1. Lipschitz continuity
If $u\in\mathcal{S}(D;N)$, then it is locally Lipschitz continuous, i.e. $u\in C^{0,1}_{\textup{loc}}(D;\mathbb{R}^{N})$, which implies that the sets